# Direct observation of the vanishing EELS cross section in graphene

### v1

@@ Abstract
Electron energy-loss spectroscopy (EELS) is a powerful tool to investigate electronic excitations in materials, as it directly probes the longitudinal dielectric response of the system, in principle resolved both in energy and momentum Egerton (2011); Ibach and Mills (1982). Special attention is often devoted to vanishing momentum transfer, also referred to as the optical limit, where the cross section contains the same information delivered by the optical response of the sample Mondio et al. (1992); Hong et al. (2021). Energy resolution of few meV can be attained, e.g., by using low-energy electrons that are reflected by the sample surface, the scattering angle enabling the measure of the momentum transfer and thus providing access to the excitation’s dispersion with a momentum resolution that has been recently pushed down to 10-3Å-1Zhu et al. (2015). This technique, known as high-resolution EELS (HREELS), is however a surface-sensitive one, and as such it cannot probe the bulk response of the target material, since the scattered electrons hardly penetrate it. On the contrary, higher energy electrons as those used in a transmission electron microscope (TEM) can pass through a target of about 1 $\mu$m thickness Egerton (2008). A sketch of the experimental geometry is shown in the left panel of Fig. 1. If the thickness $d$ of the sample is such that $q_{\mathrm{3D}}d\gg 1$, where $q_{\mathrm{3D}}=|\mathbf{k}_{i}-\mathbf{k}_{f}|$ is the momentum lost by the beam electron, the loss function can be modelled within inelastic scattering theory by approximating the dielectric response of the film with that of the bulk Sturm (1993); Egerton (2008); Keast (2016). Otherwise, finite thickness effects become relevant. Recent years have witnessed a significant effort in optimizing both energy and momentum resolution of TEM-EELS, reaching few tens of meV and few hundredths of Å-1, respectively Senga et al. (2019); Li et al. (2020); Hage et al. (2020); Senga et al. (2022); Hage et al. (2018); Hong et al. (2020); Guandalini et al. (2023a). Such advancement of the experimental setup has made TEM-EELS a valuable tool for studying bulk fundamental excitations as a function of momentum transfer in small volumes of material Pichler et al. (1998); Neudert et al. (1998); Schuster et al. (2007); Kramberger et al. (2008); Schuster et al. (2009).
@@
The angular and energy dependence of the inelastic electron scattering are specified by a double-differential cross section, that can be generally expressed as the product of two terms: the so-called loss function, accounting for the target response to an incident electron, and a kinematic prefactor that instead only depends on the scattering geometry through energy and momentum conservation Egerton (2011); Ibach and Mills (1982). The kinematic prefactor, crucial in the interpretation of HREELS spectra Mills (1975); Lambin et al. (1987); Palmer et al. (1987a); Hogan et al. (2009); Li et al. (2022) because of the dependence of the momentum transfer on both the scattering angle and the energy loss Egerton (2011), is generally simplified in TEM-EELS, where the momentum transfer is regarded to lie only within the plane perpendicular to the incoming electron beam Roth et al. (2014). This assumption is reasonable in typical experimental setups with small scattering angles and momentum resolution of few tenths of Å-1 limiting the access to the low-momentum transfer regime, and as long as the targeted film is thick enough to be viewed as a bulk 3D sample. The TEM-EELS cross section is thus proportional to $-\mathrm{Im}[\varepsilon^{-1}(q_{\mathrm{3D}},\omega)]/q_{\mathrm{3D}}^{2}$, with the energy loss $\hbar\omega$, where the loss function is given by the bulk 3D inverse dielectric function $\varepsilon^{-1}$.
@@
In the last decades, particular attention has been devoted to the study of 2D materials, i.e., materials that are periodic in two directions ($x,y$) and atomically thin in the third direction ($z$), due to their remarkable electronic and optical properties Ferrari et al. (2015); Bhimanapati et al (2015). Their intrinsic dielectric response, that in HREELS measurements is affected by the inevitable presence of substrates, can be naturally accessed by TEM-EELS enabling the study of electronic Hong et al. (2020); Woo et al. (2023) and lattice Hage et al. (2018); Senga et al. (2019) excitations of freestanding 2D materials. When the sample thickness becomes atomic and/or the system is probed near to the optical limit, dimensionality effects may arise in the dielectric response due to the partially confined nature of the electrons. In this regime, the system response becomes anisotropic, and one conveniently decomposes the momentum lost by the incoming electrons $\bm{q}_{3D}$ in $\mathbf{q}$ and $q_{z}$, the momentum transfer parallel and perpendicular, respectively, to the $xy$ plane of the system (see left panel of Fig. 1). While $\mathbf{q}$ is transferred to the crystal quasi-momentum of the 2D periodic lattice, thus being a good quantum number for the material’s electrons, $q_{z}$ is not. The latter encodes the dependence of the momentum transfer on the energy loss, while the intrinsic response of the system depends mainly on $\mathbf{q}$ Egerton (2011); Roth et al. (2014). We point out that these considerations, relevant for 2D materials, apply also to thin films, as the interpretation of EEL spectra in terms of the 3D response is an approximation valid only for $q_{\mathrm{3D}}d\gg 1$.
@@
Transmission EELS has been studied at different levels of theory mostly within a 3D framework Egerton (2011); Keast (2016); Nicholls (2021). Nonetheless, the relation between the measured EELS cross section and the dielectric response of a 2D system has been recently addressed by Nazarov Nazarov (2015), showing that it indeed differs from the 3D response. As an example, the kinematic term for the case of a 2D system at low momentum transfers significantly differs from the bulk counterpart, due to the $\approx e^{-qz}$ decay of electric fields in the vacuum. In this regime, the 2D material may be approximated as a 2D sheet Sohier et al. (2015), displaying qualitatively different dielectric properties Cudazzo et al. (2011). Remarkably, it has been theoretically predicted that the cross section of a 2D material vanishes in the low-momentum transfer regime instead of giving a maximum signal as the $1/q_{\mathrm{3D}}^{2}$ prefactor would predict, due to the inability of 2D materials to screen the Coulomb interaction in the long-wavelength limit Nazarov (2015). A TEM setup allowing for a very high momentum resolution is instrumental to probe the 2D features of the kinematic response, that we anticipate occur at $q\leq q_{z}$.
In this work, we study the EEL cross section of a prototypical 2D material, graphene, in the optical limit regime, i.e. at low-momentum transfers ($q<0.1$  Å-1). Graphene is one of the most studied 2D materials due to its potential electronic applications, e.g., for plasmonic and electronic devices Castro Neto et al. (2009); Bonaccorso et al. (2010); Grigorenko et al. (2012); García de Abajo (2014). The energy loss function of undoped graphene and of related sp2 carbon materials, graphite and nanotubes, always displays $\pi$ and $\pi+\sigma$ plasmons, arising from interband transitions occurring below $10$ eV and above $15$ eV, respectively. We demonstrate that scattering geometry effects, encoded in the kinematic prefactor are responsible for a spectral reshaping akin to HREELS Palmer et al. (1987a, b); Li et al. (2022), need to be properly disentangled from the intrinsic dielectric response of the target in order to study the relative intensities of different EEL peaks and the low-dimensional effects on screening.
@@ I Results and discussion
In the right panel of Fig. 1, we show the optical limit of the EEL cross section of graphite with a momentum resolution of $\Delta q=0.1$-1 and of freestanding graphene at the higher momentum resolution ($\Delta q=0.02$${}^{-1}\approx q_{z}$). We note the graphene cross section is completely quenched, as opposite to the case of graphite.
Figure 1: Left: parallel beam setup for momentum resolved EELS. The momentum transfer lost by the incoming electron along the direction parallel (perpendicular) to the beam, $q_{z}$ ($q$), is shown in red (blue). Right: EEL cross section at vanishing momentum transfer of graphite with standard $q$ resolution (orange) and of graphene with high resolution (blue).
As we derive in the supporting information, the electron energy-loss cross section of a 2D system near the optical limit may be written as
$\frac{d^{2}S}{d\Omega d\hbar\omega}(|\mathbf{q}|\approx 0,\omega)=\frac{4A^{u.c.}}{\pi^{2}a_{0}^{2}}\frac{q^{2}}{(q^{2}+q_{z}^{2})^{2}}\frac{\mathrm{Re}\left[\tilde{\sigma}(\omega)\right]}{\omega}\ ,$ (1)
where $a_{0}$ is the Bohr radius, $A^{\mathrm{u.c.}}$ is the area of the 2D unit cell and $\tilde{\sigma}$ is the optical conductivity along the momentum transfer $\mathbf{q}$. The EEL cross section near the optical limit contains the same information as the optical coefficients (see supporting information for more details). However, the kinematic prefactor is $I_{kin}\equiv q^{2}/(q^{2}+q_{z}^{2})^{2}$, thus the cross section is quenched for $q\leq q_{z}$, even though the optical conductivity remains finite. Using energy and momentum conservation, $q_{z}$ can be expressed asEgerton (2011); Roth et al. (2014):
@@
Figure 2: Energy loss dependence of $q_{z}$ (top panel) and the reshape factor $A$ (lower panel) for EEL spectra at different $q$ and incident energies $E_{i}$. $q_{z}$ is evaluated at $q=0$, while $A$ at $E_{i}=30$ keV. Dashed horizontal lines in the top panel mark the lowest $q$ values shown in Fig. 3. The $\pi$ and $\pi+\sigma$ plasmons of graphene lie in the grey spectral regions.
We show in Fig. 2 concrete examples of $q_{z}$ and $A$ dependencies on the energy loss $\hbar\omega$. In the top panel, we show $q_{z}$ at $q=0$ for different incoming electron energies $E_{i}$. We verified numerically that $q_{z}$ is independent on $q$ for the range considered ($q<0.1$-1, $\hbar\omega\approx$ eV). For 2D systems $q_{z}$ should be considered as a critical $q$ value under which the cross section is quenched due to 2D kinematic effects. In the lower panel of Fig. 2, we show $A$ at several $q$. The incoming electron energy $E_{i}$ is $30$ keV. Each line shows the relative cross section drop due to the increase of $q_{z}$ as a function on $\hbar\omega$. The drop in the signal is more pronounced at lower incoming energies $E_{i}$ (not shown). At sufficiently small $q$, the signal is quenched in the $\hbar\omega$ range of the valence excitations. The grey areas in Fig. 2 correspond to the energy ranges of graphene plasmons. The signal reshaping is very strong in the $\pi$ plasmon region at very low momentum transfer, while it becomes negligible at $q>0.035$Å-1. Instead the reshaping in the energy range of the $\pi+\sigma$ plasmon is approximately linear in $\omega$, and effective in the range of considered momentum transfer. $q_{z}$ and $A$ are system independent and thus apply to EEL spectra of any 2D system.
@@ I.2 Spectral reshaping at low momentum transfers
In Fig. 3, we display with blue lines a set of EEL spectra measured at different low-momentum transfers $q$ and a beam energy of $30$ keV. We emphasize the reduced width of the ZLP attained with our experimental setup, where we adopt a pinhole-type aperture instead of a slit one. We show in the right panel of Fig. 3 a sketch of the direct beam and the pinhole-type aperture at the different $\mathbf{q}$. The two lowest pinhole positions correspond to the two spectra in the bottom of Fig. 3. They are clearly different from earlier reports Kinyanjui et al. (2012); Wachsmuth et al. (2013); Nelson et al. (2014); Liou et al. (2015), but also still affected by the direct beam. The unprecedented observation of kinematically quenched plasmon peaks is attributed to the $q$ resolution of 0.02 Å-1. The pinhole-type aperture does not only block the direct beam well below 0.1 Å-1, it is also instrumental for observing the quenched plasmon peaks near the optical limit. In all the earlier reports Kinyanjui et al. (2012); Wachsmuth et al. (2013); Nelson et al. (2014); Liou et al. (2015) the loss spectra at the lowest $q$ were always the ones with the strongest plasmon peaks, provided by the spurious inclusions of electrons with $q>q_{z}$, as $\Delta q>q_{z}$.
@@
As the momentum transfer decreases down towards $0.030$-1, the $\pi+\sigma$ plasmon loses spectral weight with respect to the $\pi$ plasmon. At even lower momentum transfers, $q<0.030$-1, the spectral signatures of both the $\pi$ and $\pi+\sigma$ plasmons become comparable or smaller than the noise level, i.e., they are below the detection limit. To highlight the effect, we also plot the ×$\times 40$ zooms of the spectra for the two lowest momenta in Fig. 3. At $q\approx 0.015$-1 a small trace of the $\pi$ plasmon can still be discerned from the spectral noise but the $\pi+\sigma$ plasmon (not shown) is already indiscernible. As can be seen from the top panel of Fig. 2, $q\approx 0.015$-1 corresponds to the limit case where $q>q_{z}$ in the $\pi$ plasmon region and $q<q_{z}$ in the $\pi+\sigma$ plasmon region. At $q\approx 0.000$-1 the spectral noise is of the same order as the signal even in the energy range of the $\pi$ plasmon. This is a clear signature of the cross section quenching due to kinematic effects at very low momentum transfer. The signal quenching can be appreciated if only electrons with $q\leq q_{z}$ are collected, thus for a $q$ resolution with $\Delta q$ at least on the order of $q_{z}$ as in our experimental setup. Otherwise, spurious inclusions of electrons with $q>q_{z}$ would dominate the cross section signal. The faster quenching of the $\pi+\sigma$ spectral feature as opposed to the $\pi$ plasmon is consistent with the expected effect captured by the reshape function $A$ shown in Fig. 2. To further highlight the kinematic origin of the spectral reshaping in Fig. 3, we also plot with red lines the spectra divided by the reshape function $A$ from Eq. (4). We excluded from this procedure the $q<0.030$-1 spectra because of the too high noise enhancement. All such scaled spectra are normalized to have the same intensity as the originals at the $\pi$ plasmon peak. The removal of kinematic effects increases the $\pi+\sigma$ intensity with respect to the $\pi$ peak intensity. The decrease of the intensity ratio between the $\pi+\sigma$ and $\pi$ peaks when approaching the optical limit is thus due to kinematic effects and it is not an intrinsic property of the dielectric response of graphene.

### v3

# Direct observation of the vanishing electron energy-loss cross section in graphene

@@ Abstract
Electron energy-loss spectroscopy (EELS) is a powerful tool to investigate electronic excitations in materials, as it directly probes the longitudinal dielectric response of the system, in principle resolved both in energy and momentum [1, 2]. Special attention is often devoted to vanishing momentum transfer, also referred to as the optical limit, where the cross section contains the same information delivered by the optical response of the sample [3, 4]. Energy resolution of few meV can be attained, e.g., by using low-energy electrons that are reflected by the sample surface, the scattering angle enabling the measure of the momentum transfer and thus providing access to the excitation’s dispersion with a momentum resolution that has been recently pushed down to 10-3Å-1[5]. This technique, known as high-resolution EELS (HREELS), is however a surface-sensitive one, and as such it cannot probe the bulk response of the target material, since the scattered electrons hardly penetrate it. On the contrary, higher energy electrons as those used in a transmission electron microscope (TEM) can pass through a target of about 1 $\mu$m thickness [6]. A sketch of the experimental geometry is shown in the left panel of Fig. 1. If the thickness $d$ of the sample is such that $q_{\mathrm{3D}}d\gg 1$, where $q_{\mathrm{3D}}=|\mathbf{k}_{i}-\mathbf{k}_{f}|$ is the momentum lost by the beam electron, the loss function can be modelled within inelastic scattering theory by approximating the dielectric response of the film with that of the bulk [7, 6, 8]. Otherwise, finite thickness effects become relevant. Recent years have witnessed a significant effort in optimizing both energy and momentum resolution of TEM-EELS, reaching few tens of meV and few hundredths of Å-1, respectively [9, 10, 11, 12, 13, 14, 15]. Such advancement of the experimental setup has made TEM-EELS a valuable tool for studying bulk fundamental excitations as a function of momentum transfer in small volumes of material [16, 17, 18, 19, 20].
@@
The angular and energy dependence of the inelastic electron scattering are specified by a double-differential cross section, that can be generally expressed as the product of two terms: the so-called loss function, accounting for the target response to an incident electron, and a kinematic prefactor that instead only depends on the scattering geometry through energy and momentum conservation [1, 2]. The kinematic prefactor, crucial in the interpretation of HREELS spectra [21, 22, 23, 24, 25] because of the dependence of the momentum transfer on both the scattering angle and the energy loss [1], is generally simplified in TEM-EELS, where the momentum transfer is regarded to lie only within the plane perpendicular to the incoming electron beam [26]. This assumption is reasonable in typical experimental setups with small scattering angles and momentum resolution of few tenths of Å-1 limiting the access to the low-momentum transfer regime, and as long as the targeted film is thick enough to be viewed as a bulk 3D sample. The TEM-EELS cross section is thus proportional to $-\mathrm{Im}[\varepsilon^{-1}(q_{\mathrm{3D}},\omega)]/q_{\mathrm{3D}}^{2}$, with the energy loss $\hbar\omega$, where the loss function is given by the bulk 3D inverse dielectric function $\varepsilon^{-1}$.
@@
In the last decades, particular attention has been devoted to the study of 2D materials, i.e., materials that are periodic in two directions ($x,y$) and atomically thin in the third direction ($z$), due to their remarkable electronic and optical properties [27, 28]. Their intrinsic dielectric response, that in HREELS measurements is affected by the inevitable presence of substrates, can be naturally accessed by TEM-EELS enabling the study of electronic [14, 29] and lattice [13, 9] excitations of freestanding 2D materials. When the sample thickness becomes atomic and/or the system is probed near to the optical limit, dimensionality effects may arise in the dielectric response due to the partially confined nature of the electrons. In this regime, the system response becomes anisotropic, and one conveniently decomposes the momentum lost by the incoming electrons $\bm{q}_{3D}$ in $\mathbf{q}$ and $q_{z}$, the momentum transfer parallel and perpendicular, respectively, to the $xy$ plane of the system [see (a) panel of Fig. 1]. While $\mathbf{q}$ is transferred to the crystal quasi-momentum of the 2D periodic lattice, thus being a good quantum number for the material’s electrons, $q_{z}$ is not. The latter encodes the dependence of the momentum transfer on the energy loss, while the intrinsic response of the system depends mainly on $\mathbf{q}$ [1, 26]. We point out that these considerations, relevant for 2D materials, apply also to thin films, as the interpretation of EEL spectra in terms of the 3D response is an approximation valid only for $q_{\mathrm{3D}}d\gg 1$.
@@
Transmission EELS has been studied at different levels of theory mostly within a 3D framework [1, 8, 30]. Nonetheless, the relation between the measured EELS cross section and the dielectric response of a 2D system has been recently addressed by Nazarov [31], showing that it indeed differs from the 3D response. As an example, the kinematic term for the case of a 2D system at low momentum transfers significantly differs from the bulk counterpart, due to the $\approx e^{-qz}$ decay of electric fields in the vacuum. In this regime, the 2D material may be approximated as a 2D sheet [32], displaying qualitatively different dielectric properties [33]. Remarkably, it has been theoretically predicted that the cross section of a 2D material vanishes in the low-momentum transfer regime instead of giving a maximum signal as the $1/q_{\mathrm{3D}}^{2}$ prefactor would predict, due to the inability of 2D materials to screen the Coulomb interaction in the long-wavelength limit [31]. A TEM setup allowing for a very high momentum resolution is instrumental to probe the 2D features of the kinematic response, that we anticipate occur at $q\leq q_{z}$.
In this work, we study the EEL cross section of a prototypical 2D material, graphene, in the optical limit regime, i.e. at low-momentum transfers ($q<0.1$  Å-1). Graphene is one of the most studied 2D materials due to its potential electronic applications, e.g., for plasmonic and electronic devices [34, 35, 36, 37]. The energy loss function of undoped graphene and of related sp2 carbon materials, graphite and nanotubes, always displays $\pi$ and $\pi+\sigma$ plasmons, arising from interband transitions occurring below $10$ eV and above $15$ eV, respectively. Our graphene sample is suspended on a TEM grid and described in detail in the supporting information [38]. We consider spurious doping effects, possibly originating from the sample preparation, to be negligible, due to the lack of the Drude-plasmon contribution in the EELS. The same sample has been used in Ref. 15 to quantify the importance of excitonic effects in the position and shape of the onset and $\pi$ plasmon in EEL spectra, and we approached the optical limit as close as we could reliably analyze the gap opening after subtraction of the zero-loss peak (ZLP). Here, we move even closer to the optical limit, showing the relevance of low-momentum transfer kinematic effects, a broad topic relevant for all 2D materials which encompasses graphene physics.
We demonstrate that scattering geometry effects, encoded in the kinematic prefactor are responsible for a spectral reshaping akin to HREELS [23, 39, 25], need to be properly disentangled from the intrinsic dielectric response of the target in order to study the relative intensities of different EEL peaks and the low-dimensional effects on screening.
@@ I Results and discussion
In the (b) panel of Fig. 1, we show the optical limit of the EEL cross section of graphite with a momentum resolution of $\Delta q=0.1$-1 and of freestanding graphene at the higher momentum resolution ($\Delta q=0.02$${}^{-1}\approx q_{z}$). We note the graphene cross section is completely quenched, as opposite to the case of graphite.
Figure 1: (a): parallel beam setup for momentum resolved EELS. The momentum transfer lost by the incoming electron along the direction parallel (perpendicular) to the beam, $q_{z}$ ($q$), is shown in red (blue). (b): EEL cross section at vanishing momentum transfer of graphite with standard $q$ resolution (orange) and of graphene with high resolution (blue).
As we derive in the supporting information [38], the electron energy-loss cross section of a 2D system near the optical limit may be written as
$\frac{d^{2}S}{d\Omega d\hbar\omega}(|\mathbf{q}|\approx 0,\omega)=\frac{4}{\pi a_{0}^{2}}\frac{q^{2}}{(q^{2}+q_{z}^{2})^{2}}\frac{\mathrm{Re}\left[\tilde{\sigma}(\omega)\right]}{\omega}\ ,$ (1)
where $a_{0}$ is the Bohr radius and $\tilde{\sigma}$ is the optical conductivity along the momentum transfer $\mathbf{q}$. The EEL cross section near the optical limit contains the same information as the optical coefficients (see the supporting information [38]). However, the kinematic prefactor is $I_{kin}\equiv q^{2}/(q^{2}+q_{z}^{2})^{2}$, thus the cross section is quenched for $q\leq q_{z}$, even though the optical conductivity remains finite. Using energy and momentum conservation, $q_{z}$ can be expressed as[1, 26]:
@@
Figure 2: Energy loss dependence of $q_{z}$ (a) and the reshape factor $A$ (b) for EEL spectra at different $q$ and incident energies $E_{i}$. $q_{z}$ is evaluated at $q=0$, while $A$ at $E_{i}=30$ keV. Dashed horizontal lines in the (a) panel mark the lowest $q$ values shown in Fig. 3. The $\pi$ and $\pi+\sigma$ plasmons of graphene lie in the gray spectral regions.
We show in Fig. 2 concrete examples of $q_{z}$ and $A$ dependencies on the energy loss $\hbar\omega$. In the (a) panel, we show $q_{z}$ at $q=0$ for different incoming electron energies $E_{i}$. We verified numerically that $q_{z}$ is independent on $q$ for the range considered ($q<0.1$-1, $\hbar\omega\approx$ eV). For 2D systems $q_{z}$ should be considered as a critical $q$ value under which the cross section is quenched due to 2D kinematic effects. In the (b) panel of Fig. 2, we show $A$ at several $q$. The incoming electron energy $E_{i}$ is $30$ keV. Each line shows the relative cross section drop due to the increase of $q_{z}$ as a function on $\hbar\omega$. The drop in the signal is more pronounced at lower incoming energies $E_{i}$ (not shown). At sufficiently small $q$, the signal is quenched in the $\hbar\omega$ range of the valence excitations. The grey areas in Fig. 2 correspond to the energy ranges of graphene plasmons. The signal reshaping is very strong in the $\pi$ plasmon region at very low momentum transfer, while it becomes negligible at $q>0.035$Å-1. Instead the reshaping in the energy range of the $\pi+\sigma$ plasmon is approximately linear in $\omega$, and effective in the range of considered momentum transfer. $q_{z}$ and $A$ are system independent and thus apply to EEL spectra of any 2D system.
@@ I.2 Spectral reshaping at low momentum transfers
In Fig. 3, we display with blue lines a set of EEL spectra measured at different low-momentum transfers $q$ and a beam energy of $30$ keV. We emphasize the reduced width of the zero-loss peak (ZLP) attained with our experimental setup, where we adopt a pinhole-type aperture instead of a slit one. We show in the right panel of Fig. 3 a sketch of the direct beam and the pinhole-type aperture at the different $\mathbf{q}$. The two lowest pinhole positions correspond to the two spectra in the bottom of Fig. 3. They are clearly different from earlier reports [40, 41, 42, 43], but also still affected by the direct beam. The unprecedented observation of kinematically quenched plasmon peaks is attributed to the $q$ resolution of 0.02 Å-1. The pinhole-type aperture does not only block the direct beam well below 0.1 Å-1, it is also instrumental for observing the quenched plasmon peaks near the optical limit. In all the earlier reports [40, 41, 42, 43] the loss spectra at the lowest $q$ were always the ones with the strongest plasmon peaks, provided by the spurious inclusions of electrons with $q>q_{z}$, as $\Delta q>q_{z}$.
@@
As the momentum transfer decreases down towards $0.030$-1, the $\pi+\sigma$ plasmon loses spectral weight with respect to the $\pi$ plasmon. At even lower momentum transfers, $q<0.030$-1, the spectral signatures of both the $\pi$ and $\pi+\sigma$ plasmons become comparable or smaller than the noise level, i.e., they are below the detection limit. To highlight the effect, we also plot the ×$\times 40$ zooms of the spectra for the two lowest momenta in Fig. 3. At $q\approx 0.015$-1 a small trace of the $\pi$ plasmon can still be discerned from the spectral noise but the $\pi+\sigma$ plasmon (not shown) is already indiscernible. As can be seen from the (a) panel of Fig. 2, $q\approx 0.015$-1 corresponds to the limit case where $q>q_{z}$ in the $\pi$ plasmon region and $q<q_{z}$ in the $\pi+\sigma$ plasmon region. At $q\approx 0.000$-1 the spectral noise is of the same order as the signal even in the energy range of the $\pi$ plasmon. This is a clear signature of the cross section quenching due to kinematic effects at very low momentum transfer. The signal quenching can be appreciated if only electrons with $q\leq q_{z}$ are collected, thus for a $q$ resolution with $\Delta q$ at least on the order of $q_{z}$ as in our experimental setup. Otherwise, spurious inclusions of electrons with $q>q_{z}$ would dominate the cross section signal. The faster quenching of the $\pi+\sigma$ spectral feature as opposed to the $\pi$ plasmon is consistent with the expected effect captured by the reshape function $A$ shown in Fig. 2. To further highlight the kinematic origin of the spectral reshaping in Fig. 3, we also plot with red lines the spectra divided by the reshape function $A$ from Eq. (4). We excluded from this procedure the $q<0.030$-1 spectra because of the too high noise enhancement. All such scaled spectra are normalized to have the same intensity as the originals at the $\pi$ plasmon peak. The removal of kinematic effects increases the $\pi+\sigma$ intensity with respect to the $\pi$ peak intensity. The decrease of the intensity ratio between the $\pi+\sigma$ and $\pi$ peaks when approaching the optical limit is thus due to kinematic effects and it is not an intrinsic property of the dielectric response of graphene.